Park hotel is franchised by a company that sets a base franchisee fee as $7,200 and gets $14.00 from every room sold. If Park hotel has 400 rooms with 70.00% occupancy for December, what is the unit mixed cost (UMC) for Park in December? Group of answer choices $28.17 $53.86 $39.71 $59.32
Answer #1

frenchise fees- 7200

sale of rooms- 400 X 70% X 14= 3920

total cost= 7200+3920= 11120

unit mixed cost= Total cost/ Number of rooms

= 11120/280= 39.71

So option 3 is correct